OpenStudy (baconfuntimes):

What is a+b/q when q equals 7 and b equals 17 and a equals 15

OpenStudy (texaschic101):

Is there any parenthesis in this problem ?

OpenStudy (phi):

is this \[ \frac{a+b}{q}\] or \[ a + \frac{b}{q} \]

OpenStudy (baconfuntimes):

first one

OpenStudy (phi):

can you replace the letters with the numbers they give you ?

OpenStudy (texaschic101):

this is just a matter of substitution and solving....not really that hard

OpenStudy (baconfuntimes):

yes 17+15/7 and there is remainder

OpenStudy (baconfuntimes):

oh it 4r4

OpenStudy (texaschic101):

leave it a fraction

OpenStudy (phi):

ok, but you should put parens around the first two numbers (17+15)/7 because you are really doing \[ \frac{17+15}{7}\]

OpenStudy (baconfuntimes):

oh

OpenStudy (phi):

you can write the answer a few ways. If this is multiple choice, then write it to match one of the choices. otherwise you can keep it an improper fraction or write it as a mixed number 4 4/7 (that is 4 and 4/7)

OpenStudy (texaschic101):

or 32/7

OpenStudy (baconfuntimes):

this is how the equation is show

OpenStudy (texaschic101):

but your resulting answer can be written in more then 1 way...it can be left as an improper fraction of 32/7....or it can be changed to a mixed number 4 4/7

OpenStudy (phi):

yes, and that means (a+b)/q which can also be written as \[ \frac{a}{q}+\frac{b}{q}\] that might be confusing, but it is the same as adding two fractions with the same denominator either way you will get 4 and 4/7

OpenStudy (baconfuntimes):

ok thanks